### v1

# Boter: Bootstrapping Knowledge Selection and Question Answering for Knowledge-based VQA

Abstract.
Knowledge-based Visual Question Answering (VQA) requires models to incorporate external knowledge to respond to questions about visual content. Previous methods mostly follow the ”retrieve and generate” paradigm. Initially, they utilize a pre-trained retriever to fetch relevant knowledge documents, subsequently employing them to generate answers. While these methods have demonstrated commendable performance in the task, they possess limitations: (1) they employ an independent retriever to acquire knowledge solely based on the similarity between the query and knowledge embeddings, without assessing whether the knowledge document is truly conducive to helping answer the question; (2) they convert the image into text and then conduct retrieval and answering in natural language space, which may not ensure comprehensive acquisition of all image information. To address these limitations, we propose Boter, a novel framework designed to bootstrap knowledge selection and question answering by leveraging the robust multimodal perception capabilities of the Multimodal Large Language Model (MLLM). The framework consists of two modules: Selector and Answerer, where both are initialized by the MLLM and parameter-efficiently finetuned in a simple cycle: find key knowledge in the retrieved knowledge documents using the Selector, and then use them to finetune the Answerer to predict answers; obtain the pseudo-labels of key knowledge documents based on the predictions of the Answerer and weak supervision labels, and then finetune the Selector to select key knowledge; repeat. Our framework significantly enhances the performance of the baseline on the challenging open-domain Knowledge-based VQA benchmark, OK-VQA, achieving a state-of-the-art accuracy of 62.83%. We conduct comprehensive ablations of our framework, including Selector and Answerer ablations, the method of cycle training, varying key knowledge document numbers, the impact of vision information, and so on.
Knowledge-based VQA, Selector and Answerer, Multimodal Large Language Model, Cycle training
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06
Figure 1. Comparison with previous paradigms. (a) Previous KB-VQA paradigms apply captioning tools to transform the image into a visual context. They use a separate pre-trained knowledge retriever to retrieve knowledge and then use them to predict answers in natural language space. However, the top-ranked document contains noise information and misleads the model to answer a wrong answer ”surfing”. (b) The proposed paradigm aims to enhance the utilization of visual information to perform knowledge selection and question-answering consistency. We bootstrap the selector module and answerer module each other in a cycle style: where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector. The selector successfully finds key knowledge documents that are low-ranked but informative and the answerer predicts the right answer ”freestyle”.
1. Introduction
Visual Question Answering (VQA) (Antol et al., 2015) is a task that entails answering questions based on an image. Conventional VQA benchmarks (Goyal et al., 2017; Hudson and Manning, 2019) primarily focus on addressing straightforward questions that only necessitate visual perception and recognition. However, in real-world scenarios, many questions require broader world knowledge and common sense for accurate responses. Consequently, the Knowledge-based VQA (KB-VQA) benchmark is introduced to facilitate the development of models suitable for real-world applications. In contrast to previous VQA tasks, KB-VQA presents a more challenging setting where questions grounded in visual content demand answers by referencing external knowledge. Early studies (Wang et al., 2017a, b) offer a pre-defined knowledge base for all questions. More recently, the Outside-Knowledge Visual Question Answering (OK-VQA) (Marino et al., 2019) benchmark is introduced to emphasize the utilization of open-domain world knowledge to answer questions, rather than relying on a predefined knowledge base. This approach simulates a more realistic and human-like interaction with the visual world.
To tackle the knowledge-based Visual Question Answering (VQA) task, a common approach involves retrieving relevant explicit knowledge and leveraging it to answer questions. Existing KB-VQA methods (Marino et al., 2021; Gardères et al., 2020; Lin and Byrne, 2022; Gui et al., 2021; Lin et al., 2022, 2023) often utilize the pre-trained retriever (e.g., DPR (Karpukhin et al., 2020)) to retrieve the related knowledge from diverse explicit knowledge bases, such as ConceptNet (Speer et al., 2017), Wikipedia (Vrandečić and Krötzsch, 2014), and Google Search Corpus (Luo et al., 2021). Subsequently, they utilize the retrieved knowledge to generate answers based on pre-trained language models. In addition to methods utilizing explicit knowledge, some approaches (Gui et al., 2021; Lin et al., 2022; Hu et al., 2022a; Shao et al., 2023) employ GPT-3 (Brown et al., 2020), a large-scale pre-trained language model, to generate implicit textual knowledge relevant to the question. With the explicit and implicit knowledge, the models have achieved good results. In addition to augmenting various knowledge bases, two critical components should be considered to obtain good performance: vision information mining and knowledge selection.
To obtain vision information, many traditional methods transform the image into raw texts, such as captions (Hu et al., 2022a; Lin et al., 2022; Gui et al., 2021), tags (Yang et al., 2022), and OCR (Gao et al., 2022; Lin and Byrne, 2022). They use textual visual information in the process of retrieving and answering, which leads to the underutilization of visual information. On the other hand, to obtain external knowledge, many methods simply use a pre-trained retriever to obtain the relevant knowledge, without considering how to make the answer selector work better with the answer generator. The retrieved documents are only ranked based on the similarity between query and knowledge embeddings, which cannot guarantee that the selected documents can genuinely aid in producing correct answers.
To address the aforementioned limitations, we introduce a novel framework, Boter, designed to select key knowledge documents and answer questions using the Multimodal Large Language Model (MLLM). We aim to enhance the utilization of visual information by leveraging the MLLM’s multimodal perception capabilities.
Built upon the MLLM, our framework comprises two modules: a Selector and an Answerer. We train two modules by repeating the following process: the Selector first identifies important knowledge from the candidate knowledge documents retrieved by the pre-trained retriever; then, the Answerer takes the key knowledge documents as the input knowledge and is finetuned to generate the answer; next, we generate pseudo-labels of key knowledge documents according to the Answerer’s predictions and weak supervision labels; finally, we refine the Selector to assess the relevance of retrieved knowledge documents in answering the question. This strategy of cycle training bootstraps the ability of knowledge selection and answer generation consistently, enabling the model to accurately respond to knowledge-intensive questions.
We conduct extensive experiments on the open-domain knowledge-based VQA benchmark (OK-VQA (Marino et al., 2019)) to validate the effectiveness of the proposed framework, where our method largely outperforms the baseline and achieves the state-of-the-art performance of 62.83%, only finetuning 0.16% parameters with LoRA (Hu et al., 2022b). We also conduct comprehensive ablations to validate the impact of different components of the proposed framework, including the Effect of Selector and Answerer, cycle training of the framework, varying the number of key knowledge documents, the impact of vision information, and so on.
Our contributions are summarized as follows:
  • We introduce a novel framework for consistently selecting key knowledge documents and answering questions using the multimodal language model.
  • We propose a new cycle-training method to bootstrap Selector and Answerer each other, where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector.
  • We achieve a state-of-the-art performance of 62.83% on the OK-VQA dataset, surpassing the previous state-of-the-art method. Notably, this improvement is achieved by fine-tuning only 0.16
  • We present a comprehensive analysis to assess the impact of various components within the proposed framework.
2. RELATED WORK
Visual Question Answering (VQA). VQA presents a formidable challenge in visual-language understanding, requiring models to respond to visually-grounded questions. Conventional VQA benchmarks (Goyal et al., 2017; Hudson and Manning, 2019) primarily focus on basic visual perception and reasoning tasks. Numerous studies have achieved promising results on these benchmarks by incorporating: (1) improved visual features (Anderson et al., 2017; Jiang et al., 2020; Shen et al., 2021; Zhang et al., 2021); (2) attention-based cross-modality fusion techniques (Hu et al., 2017; Kim et al., 2018; Li et al., 2019; Yu et al., 2019); (3) the pretrain-finetune paradigm (Tan and Bansal, 2019; Lu et al., 2019; Zhang et al., 2021; Li et al., 2022, 2020; Wang et al., 2022). In addition to these general-purpose VQA tasks, the knowledge-based VQA task (Wang et al., 2017a; Marino et al., 2019) is introduced to advance models’ understanding of the open-ended world. This task leverages world knowledge to answer visual questions, thus enabling models to grasp the broader context of visual queries.
Figure 2. Our Boter framework consists of two modules: a Selector and an Answerer. Selector (left) selects the top-T knowledge documents for the Answerer (right), and the Answerer focuses on important knowledge information to predict answers. Both modules utilize the same frozen visual module to extract image features. We train the fully connected (FC) layer and fine-tune the language model using LoRA, which amounts to only 0.16% of the total parameters. For detailed training procedures of the two modules, refer to Alg. 1. The original knowledge is retrieved using DPR, and for brevity, we omit the retrieval process here (details can be found in Section 3.1).
Knowledge-based VQA. Early benchmarks (Wang et al., 2017a, 2015) provide fixed knowledge bases for the model to retrieve and answer questions. Recently, the open-domain knowledge-based VQA benchmark (OK-VQA (Marino et al., 2019)) is proposed, which provides no extra knowledge bases and the model needs to find proper knowledge bases to help answer questions. Recent studies (Gardères et al., 2020; Wu et al., 2022; Lin and Byrne, 2022; Gui et al., 2021) have explored various open-domain world knowledge sources, such as ConceptNet (Speer et al., 2017), Wikipedia (Vrandečić and Krötzsch, 2014), Google Search Corpus (Luo et al., 2021). They retrieve the relevant knowledge documents from the knowledge bases and integrate them into the answering model to generate predictions. Except for using explicit knowledge, some methods also take GPT-3 (Brown et al., 2020) as an implicit knowledge producer. They either prompt GPT-3 with in-context examples to predict answers directly (Yang et al., 2022; Hu et al., 2022a; Shao et al., 2023), or use GPT-3 to generate answer candidates with evidence serving as textual implicit knowledge bases (Gui et al., 2021; Lin et al., 2022), leading to significant performance improvements.
Different from previous approaches that primarily enhance performance by incorporating more knowledge sources, our focus lies in effectively leveraging visual information and retrieved knowledge. We employ a multimodal large language model to enhance visual understanding and construct a Selector and an Answerer to consistently select key knowledge and answer questions.
Multimodal large language model. Recently, multimodal large language models (Li et al., 2023; Alayrac et al., 2022; Liu et al., 2023; Dai et al., 2023) have demonstrated remarkable visual-language understanding and reasoning capabilities, owing to the advancement of larger language models (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023; vic, 2023; Touvron et al., 2023). These methods typically consist of a frozen visual encoder (Radford et al., 2021), a visual-language connector (Li et al., 2023), and a large language model (Chung et al., 2022; Zhang et al., 2022; vic, 2023). The models are firstly pre-trained on large-scale visual-text datasets to align visual features to the language embedding space. After pretraining, the large language model can understand the visual details. Then, the model is finetuned to adapt to various visual-language tasks. In this study, we adopt BLIP2, one of the widely used models, as our backbone for understanding visual details and bootstrap knowledge selection and question answering with it.
3. Method
In this section, we first introduce the preliminaries of Knowledge Retrieval and MLLM, which are the foundation of our framework. Then, we present the design of the Selector and Answerer for knowledge selection and question answering on knowledge respectively. Finally, we illustrate the cycle training method of two designed modules.
3.1. Preliminaries
Knowledge Retrieval. Following previous knowledge-based VQA methods (Gao et al., 2022; Lin and Byrne, 2022; Si et al., 2023), we adopt the dense passage retrieval (DPR) (Karpukhin et al., 2020) to retrieve the knowledge documents. We transform the visual image $I_{i}$ into texts through the following steps:
  • The pretrained captioning model Oscar+ (Zhang et al., 2021) is adopted to obtain the image caption $C_{i}$.
  • The pretrained detection model VinVL (Zhang et al., 2021) is adopted to extract the text description of objects and attributes $L_{i}$.
  • The Google OCR (Optical Character Recognition) APIs are adopted to extract the text strings $\{O_{i}\}$ of the image.
Consequently, the visual contexts of the image $I$ can be represented as $V_{i}=(C_{i},L_{i},O_{i})$. We combine the question $Q_{i}$ and the visual contexts $V_{i}$ as a query $q_{i}=(Q_{i},V_{i})$ to retrieve the relevant documents from the knowledge base $D$. DPR encodes the query $q_{i}$ and all the knowledge documents $\mathcal{D}$ into embeddings $\mathbf{q}_{i}$ and $\{\mathbf{d}_{1},\mathbf{d}_{2},...,\mathbf{d}_{n}\}$ with different BERT-like (Devlin et al., 2018) encoders as follows:
(1) $\mathbf{q}_{i}=BERT_{q}(q_{i}),\mathbf{d}_{i}=BERT_{d}(D_{i}),$
We compute the similarity scores $sim(q_{i},D_{j})=\mathbf{q}_{i}^{T}\cdot\mathbf{d}_{j}$ and exploit FAISS (Johnson et al., 2019) to index Top-k related knowledge documents $\mathcal{P}_{i}=\{P_{i,1},P_{i,2},...,P_{i,k}\}$ for $i$-th query.
Multimodal Large Language Model. Recently, multimodal large language models (Li et al., 2023; Alayrac et al., 2022; Liu et al., 2023) have demonstrated significant advancements in visual-language understanding. The architecture of these models typically comprises a frozen image encoder (Dosovitskiy et al., 2020; Fang et al., 2023), a visual-language connector (Li et al., 2023), and a pre-trained language model (Chung et al., 2022; Zhang et al., 2022; vic, 2023). Given an image $I_{i}$, the frozen image encoder outputs a set of visual features $\{\mathbf{h}_{i,1},\mathbf{h}_{i,2},...,\mathbf{h}_{i,m}\}$. Visual-language connector takes extracted visual features as input, and outputs language-aligned visual features $\{\mathbf{v}_{i,1},\mathbf{v}_{i,2},...,\mathbf{v}_{i,l}\}$. These visual features are concatenated with the textual word embeddings, which are fed into the language model for generation. Through pre-training on large-scale image-caption datasets, the visual-language connector can effectively project visual features into the feature space of the Language Large Model (LLM).
In our work, both knowledge selection and question-answering modules adopt BLIP-2 (Li et al., 2023) as the backbone. We freeze the visual encoder and Q-former during training. We train the fully connected layer and use LoRA (Hu et al., 2022b) to finetune the LLM (only finetune 0.16% of total parameters).
3.2. Selector and Answerer
For the knowledge-based VQA task, previous methods (Lin and Byrne, 2022; Gui et al., 2021; Lin et al., 2022) separately use DPR to retrieve relevant knowledge documents from the knowledge base, and then directly use them to help produce the answer for the given question. While these methods have demonstrated the effectiveness of leveraging external knowledge, the separate retriever is not inherently aligned with the answerer. The retrieved knowledge documents are ranked based on similarity scores between the query and knowledge embeddings, which may not always be useful for answering a specific question. Additionally, these methods transform the image into textual information and conduct retrieval and answering in the language space, resulting in suboptimal utilization of visual information.
We adopt a multimodal large language model, BLIP-2 (Li et al., 2023), to enhance the utilization of visual information. We assign BLIP-2 two roles: a Selector and an Answerer. The Selector is responsible for selecting key knowledge from the retrieved documents that truly assist in answering questions, while the Answerer predicts the answer based on the selected knowledge. This integrated approach ensures better alignment between knowledge selection and question answering and enhances the model’s ability to answer questions.
Selector. After obtaining the Top-k knowledge documents using DPR for the $i$-th sample, we aim to choose $t$ most important knowledge documents from the retrieved documents. where $t$ is smaller than $k$. As shown in Fig. 2, we firstly use the frozen image encoder and Q-former to extract the image features $\mathbf{V}_{i}$, where these features are extracted once and then used by the Selector and the Answerer. Then image features $\mathbf{V}_{i}$ are fed into the independent fully-connected layer to obtain the visual embeddings $\mathbf{E}^{v}_{i}$. We concatenate the question, a retrieved knowledge document, and the Selection prompt ”Does the retrieved knowledge document provide the key information to help answer the question?” into one sentence $S$. Next, visual embeddings $\mathbf{E}^{v}_{i}$ and the text are concatenated and fed into the LLM (Flan-T5 (Chung et al., 2022) is adopted in our work). Last, we use the probability of generating the word ‘yes’ as the score of each retrieved knowledge document $P_{i,j}$, denoted as $s_{i,j}=\bm{LLM}(concat(\mathbf{E}^{v}_{i},S_{i}))$, and we select top-t documents $\hat{\mathcal{P}}_{i}=\{\hat{P}_{i,1},\hat{P}_{i,2},...,\hat{P}_{i,t}\}$ based on the scores. The Selector can be conceptualized as follows:
(2) $\hat{\mathcal{P}}_{i}==Selector(I_{i},Q_{i},\mathcal{P}_{i}),|\hat{\mathcal{P}}_{i}|=t$
Answerer. After obtaining the selected knowledge documents, we aim to reason on the knowledge to answer questions. As shown in Fig. 2, we process the same image features to obtain the different visual embeddings $\mathbf{E}^{v}_{i}$ via the fully-connected layer of the Answerer. Next, we concatenate the question and the knowledge into one sentence $S^{\prime}$ using the template ”Question: {} Knowledge: {} Answer: ”. We concatenate the visual embeddings and the text, which are fed into the LLM with different LoRA parameters to get the answer. The model outputs corresponding answers based on different documents. The Answerer can be conceptualized as follows:
(3) $a_{i}=Answerer(I_{i},Q_{i},\hat{\mathcal{P}}_{i})$
Then the final answer is based on the majority vote. We also tried different knowledge reasoning methods, such as concatenating (the results can be seen in the ablation study).
Algorithm 1 Pipeline of cycle training
  Input: KB-VQA dataset $\mathcal{D}=\{I_{i},Q_{i},\mathcal{A}_{i}|i=1,2,\dots,N\}$; Retrieved knowledge documents $\mathcal{P}_{i}=\{P_{i}^{1},P_{i}^{2},\dots,P_{i}^{k}\}$; $I_{i}$, $Q_{i}$, $\mathcal{P}_{i}$, and $\mathcal{A}_{i}$ denote image, question, document set, and answer set of $i$-th sample
  Output: Knowledge selection model $Selector$; Question answering model $Answerer$
  for sample in $\mathcal{D}$ do
     Stage 1:
     1: Using $Selector$ to select top-t documents $\mathcal{\hat{P}}_{i}$ from the retrieved knowledge documents $\mathcal{P}_{i}$ as Eq. 2
     2: Finetuning $Answerer$ on $\{I_{i},Q_{i},\mathcal{\hat{P}}_{i},\mathcal{A}_{i}\}$ supervised by the ground-truth answer as Eq. 4.
     Stage 2:
     1: Using $Answerer$ to predict answers for retrieved knowledge documents $\mathcal{P}_{i}$ as Eq. 3
     2: Generating to pseudo labels $\{y_{i,j}\}$ for retrieved knowledge documents $\mathcal{P}_{i}$ as Eq. 5
     3: Finetuning $Selector$ on $\{I_{i},Q_{i},\mathcal{P}_{i},\{y_{i,j}\}\}$ supervised by the pseudo label as Eq. 6.
  end for
3.3. Cycle Training
To enable the Selector and Answerer to select key knowledge and answer questions, we bootstrap them with each other in a style of cycle training. We repeat the following process for the given $i$-th sample $\{I_{i},Q_{i},\mathcal{P}_{i},\mathcal{A}_{i}\}$ of the training dataset:
Answerer Training. We use Eq. 2 to get the selected knowledge documents $\hat{\mathcal{P}}_{i}$. The image $I_{i}$ is fed into the frozen ViT and Q-former to obtain the image features $\mathbf{V}_{i}$. We use the trainable $FC_{ans}$ layer to output the visual embeddings $\mathbf{E}^{v}_{ans,i}$. We concatenate the visual embedding, the question $Q_{i}$ and each selected knowledge document $\hat{P}_{i,j}$ to construct $t$ triplets for the sample, where $j={1,2,\dots,t}$. Then we finetune the Answerer with LoRA under the supervision of the ground truth answer $\mathcal{A}_{i}$:
(4) $\mathbf{E}^{v}_{ans,i}=FC_{ans}(\mathbf{V}_{i}),$
$L_{ans}=-\sum_{j=1}^{t}\log LLM_{ans}(a_{i}^{*}|\mathbf{E}^{v}_{ans,i},Q_{i},\hat{P}_{i}^{j}),$
where $a_{i}^{*}$ is the most frequent answer in the human-annotated answer set $\mathcal{A}_{i}$.
Selector Training. We first use Eq. 3 to predict answers based on each retrieved knowledge document $P_{i,j}$. Then we assign pseudo labels to the retrieved documents according to model predictions and weak supervision labels (Luo et al., 2021; Lin and Byrne, 2022; Lin et al., 2023). We use ”yes” and ”no” as pseudo labels, where label a document as positive knowledge if Answerer can output the correct answer using that document and the document contains any of the answers in $\mathcal{A}_{i}$.
(5) $y_{i,j}=\begin{cases}\text{yes},\quad\textit{if}\ \ a_{i}=a_{i}^{*}\wedge P_{i,j}\ \ \text{contains an answer in}\ \ \mathcal{A}_{i}\\ \text{no},\quad else\end{cases}$
Table 1. Performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset. Knowledge Sources: ConceptNet (C); Wikipedia (W); Google Search (GS); Google Images (GI). The best result in the table is bolded. Compared with the previous SOTA method Prophet (Shao et al., 2023) using GPT-3, our method achieves 62.83% performance, with a performance improvement of 1.73%. The results demonstrate the effectiveness of our method.
Models Large Models $K_{train}$ $K_{test}$ Knowledge Resource Accuracy (%)
BAN+AN (Marino et al., 2019) - - - W 25.6
ConceptBERT (Gardères et al., 2020) - - - C 33.7
KRISP (Marino et al., 2021) - - - C+W 38.4
Visual Retriever-Reader (Luo et al., 2021) - 100 100 GS 39.2
MAVEx (Wu et al., 2022) - - - W+C + GI 39.4
TRiG(Ensemble) (Gao et al., 2022) T5-large (770M) 100 100 W 50.5
KAT(Single) (Gui et al., 2021) T5-large (770M) 40 40 W + GPT-3 (175B) 53.1
KAT(Ensemble) (Gui et al., 2021) T5-large (770M) 40 40 W + GPT-3 (175B) 54.4
RA-VQA (Lin and Byrne, 2022) T5-large (770M) 5 50 GS 54.5
REVIVE(Single) (Lin et al., 2022) T5-large (770M) 40 40 W+GPT-3 (175B) 56.6
REVIVE(Ensemble) (Lin et al., 2022) T5-large (770M) 40 40 W+GPT-3 (175B) 58.0
BLIP2 T5-XXL (Dai et al., 2023) BLIP2 T5-XXL (11B) - - BLIP2 54.7
PICa (Yang et al., 2022) GPT-3 (175B) - - GPT-3 (175B) 48.0
Flamingo (Alayrac et al., 2022) Flamingo (80B) - - Flamingo (80B) 57.8
PromptCap (Hu et al., 2022a) GPT-3 (175B) - - GPT-3 (175B) 60.4
Prophet (Shao et al., 2023) GPT-3 (175B) - - GPT-3 (175B)+MCAN 61.1
Boter(Ours) BLIP2 T5-XL (3B) 5 5 GS 62.83
After obtaining the pseudo label of each retrieved knowledge document, we use the trainable $FC_{sel}$ layer to output the visual embeddings $\mathbf{E}^{v}_{sel,i}$. we concatenate the visual embedding, the question $Q_{i}$ and each retrieved knowledge document $P_{i,j}$ to construct $k$ triplets for the sample, where $j={1,2,\dots,k}$. Then we finetune the Selector with LoRA under the supervision of pseudo labels:
(6) $\mathbf{E}^{v}_{sel,i}=FC_{sel}(\mathbf{V}_{i}),$
$L_{sel}=-\sum_{j=1}^{k}\log LLM_{sel}(y_{i,j}|\mathbf{E}^{v}_{sel,i},Q_{i},P_{i}^{j})$
We provide the overall training pipeline in Alg. 1. Through continuous iteration, the Selector will provide more crucial knowledge for the Answerer to accurately respond to questions. Meanwhile, the improvement in the Answerer’s reasoning ability will also result in more precise pseudo-labeling, further enhancing the Selector’s discriminative power. During the inference stage, we utilize the Selector to choose key knowledge, and then instruct the Answerer to respond to questions based on this knowledge.
4. EXPERIMENTS
4.1. Experimental Setup
Dataset. We conducted extensive experiments on OK-VQA (Marino et al., 2019) to evaluate the effectiveness of our method. OK-VQA is a challenging open-domain knowledge-based VQA dataset that requires models to leverage various external knowledge sources to answer questions. The dataset contains 14,055 questions and 14,031 images, whereas the training set and testing set have 9k and 5k image-question pairs, respectively. Due to no knowledge base being provided for OK-VQA, we need to choose the proper knowledge base for the dataset. In this paper, we adopt Google Search Corpus (Luo et al., 2021) as the knowledge base which is collected in the websites using the Google Search API. The dataset has been proven to be effective by previous methods (Luo et al., 2021; Lin and Byrne, 2022; Lin et al., 2023). We do not evaluate our method on A-OKVQA (Schwenk et al., 2022) like other retrieved-based methods (Gui et al., 2021; Lin et al., 2022; Lin and Byrne, 2022; Lin et al., 2023), as it emphasizes visually-grounded reasoning rather than knowledge retrieval and selection.
Evaluation Metric. We use the standard VQA metric (Antol et al., 2015) to evaluate the performance of the model. Given the prediction of the question $a$ and the groudtruth answer set $\mathcal{A}$, the VQA accuracy is calculated as:
(7) $Accuracy(a,\mathcal{A})=\min(\frac{\#A(a)}{3},1),$
where the groudtruth answer set $\mathcal{A}$ is annotated by different humans, #$\#A(a)$ denotes the occurrence of $a$ in $\mathcal{A}$.
Implementation Details. In our experiment, we adopt BLIP2 T5-XL (3B) (Li et al., 2023) to initialize the Selector and Retriever. We freeze the image encoder and Q-former, with both the Selector and Retriever sharing the same visual module. We finetune the fully connected layer and use LoRA (Hu et al., 2022b) to train the LLM. We use the default huggingface-PEFT setting: r=8, lora_alpha=32, lora_dropout=0.1. We use Adam as the optimizer and set the batch size to 8. We use the warm-up strategy which trains the model with an initial learning rate of 1e-4 and warm-up factor of 0.05 for 1000 steps and then utilizes a cosine annealing learning strategy with an initial learning rate of 1e-4 and a final learning rate of 0 after 10 epochs. We use top-30 knowledge documents retrieved by a pre-trained DPR (Karpukhin et al., 2020) as candidates for Selector and use the selected top-5 documents from the 30 documents for the Answerer to train and infer, denoted as $K_{candidate}=30,K_{train}=5,K_{test}=5$. We use 2 Nvidia A800 GPUs (80G) for all experiments. And our codes will be released upon paper acceptance.
4.2. Overall Results
We demonstrate the performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset in the Tab. 1. From the results, we have the following observations: (1) Early models (BAN+AN (Marino et al., 2019), ConceptBERT (Gardères et al., 2020), KRISP (Marino et al., 2021), Visual Retriever-Reader (Luo et al., 2021), and MAVEx (Wu et al., 2022)) adopt the small visual-language model as backbones and inject various knowledge bases (ConceptNet, Wikipedia, Google Web Search and Google Images) into the model, achieving a VQA accuracy from 25.6% to 39.4%. Although these models use external knowledge to help produce answers, they have relatively low performance because of the weak knowledge reasoning ability of traditional backbones. (2) Recent methods (TRiG(Gao et al., 2022), KAT (Gui et al., 2021), RA-VQA (Lin and Byrne, 2022), and REVIVE (Lin et al., 2022)) introduce the large language model to improve the capacity for common-sense reasoning. They exploit T5-large (Chung et al., 2022) to reason on the knowledge for question answering. Compared with early methods, they have a higher accuracy ranging from 50.5% to 58.0%. (3) Some recent methods take a very large language model GPT-3 (Brown et al., 2020) as the implicit knowledge base for OK-VQA. PICa (Yang et al., 2022) is the first method to use GPT-3. It uses the style of in-context learning to prompt GPT-3 to answer questions directly. Inspired by PICa, some methods (Hu et al., 2022a; Shao et al., 2023) improve the performance of GPT-3 by generating question-guided captions and providing answer candidates, achieving higher VQA score on OK-VQA, about 61%. Meanwhile, some retrieved-based methods (Gui et al., 2021; Lin et al., 2022) use GPT-3 to generate implicit knowledge and combine them with the retrieved explicit knowledge, which improves the performance. (4) Recent large multimodal large language models have achieved high performance on OK-VQA without using external knowledge (e.g., Flamingo (Alayrac et al., 2022) (80B) has a zero-shot VQA score of 57.8%, BLIP2 T5-XXL (Dai et al., 2023) has a finetuning VQA score of 54.7%). (5) Our method aims to select key knowledge and answer questions utilizing the multimodal perception and reasoning ability of the multimodal large language model. By cycle training, we bootstrap Selector and Answerer to perform knowledge selection and question answering consistently. With a small number of selected knowledge documents, our method outperforms other state-of-the-art models, achieving an accuracy of 62.83%. It is worth noting that we do not use GPT-3 and we only train the 0.16% parameters of the model.
4.3. Ablation Study
We conduct the ablation studies to evaluate different components of our framework on OK-VQA.
Effect of Selector. We conduct the ablation study to evaluate the effectiveness of Selector in our method. We show the results in Tab. 2. From the results, we can observe: (1) Our framework, leveraging key knowledge documents selected by the Selector, consistently outperforms the Answerer when using the same number of documents retrieved by DPR. We improve the performance by 2.14% and 1.88% with 1 and 5 test knowledge documents, compared to DPR-based retrieval. When using the randomly selected documents, the model performs worst. These results demonstrate that top-ranked knowledge documents based on DPR scores are not optimal for question answering and our key knowledge selection module can identify relevant documents for accurate question answering, ensuring the coherence of knowledge retrieval and question-answering processes. (2) With an increase in the number of test knowledge documents to 10, the performance of the model DPR-based retrieval almost has no improvement and our method’s performance decreases. This suggests that additional documents may introduce more irrelevant knowledge, adversely affecting question-answering accuracy.
Table 2. Ablation study on the Selector. We select 5 knowledge documents from top-30 knowledge candidates retrieved by DPR. ’DPR Score’ refers to selecting top-5 knowledge based on similarity scores. ’Random Selection’ means randomly selecting 5 knowledge documents from 30 candidate knowledge documents. ’Selector’ denotes choosing 5 key knowledge documents by the Selector.
$K_{train}$ $K_{test}$ Knowledge Selection Accuracy (%)
5 1 Random Selection 50.45
5 1 DPR Score 58.80
5 1 Selector 61.62
5 5 Random Selection 55.05
5 5 DPR Score 60.69
5 5 Selector 62.83
5 10 Random Selection 52.11
5 10 DPR Score 60.93
5 10 Selector 62.17
Effect of Answerer. In Tab. 3, we present a comparison of our Answerer using different knowledge reasoning methods. The results show that the performance using the strategy of voting surpasses that of concatenating under different knowledge selection settings. We argue that directly combining all the knowledge documents into a lengthened document makes it difficult for the Answerer to reason on them, which is easily influenced by noisy information. In contrast, it is easier for the Answerer to reason on each document to predict the answer. Simple voting can choose the best answer.
Table 3. Ablation study on different knowledge reasoning methods of Answerer. ’Concatenating’ denotes that we combine the key knowledge documents into one sentence and feed it into the Answerer to predict the final answer. ’Voting’ means that we feed different key knowledge documents into the Answerer to predict different answers and choose the best answer based on majority voting.
Method Backbone Accuracy (%)
Concatenating BLIP2 (fine-tuned) w knowledge from DPR 59.11
Voting 60.69
Concatenating Boter (Ours) w knowledge from Selector 62.06
Voting 62.83
Effect of cycle training. To evaluate the effectiveness of our cycle-training method, we compare the method with the strategy of independent training of two modules. We finetune the Answerer with the knowledge documents retrieved by DPR as the baseline. Independent training means that we train the Selector and the Answerer respectively. Initially, we train the Answerer module utilizing knowledge documents retrieved by DPR. Subsequently, employing the trained Answerer, we generate answers for each retrieved knowledge document, thereby generating pseudo-labels for the retrieved knowledge. We then proceed to train the Selector module supervised by these pseudo-labels. Finally, we conduct finetuning of the Answerer once more, incorporating new knowledge documents selected by the Selector. The results in Tab. 4 show that the model with cycle training outperforms the model with iterative refinement by 3.81%. The VQA score of using independent training is even lower than the baseline. These results demonstrate that our cycle training method can effectively boost the Selector and Answerer each other, which makes the model find key knowledge documents and leverage the knowledge to answer questions.
Table 4. Ablation study on different training methods of our framework.
Methods Accuracy (%)
Baseline 60.69
Independent training 59.02
Cycle training 62.83
Effect of different methods of pseudo-labeling. In Tab. 5, we compare the model performance with different methods of pseudo-labeling. When using the model predictions as guidance, the model has a VQA score of 62.31%. When adding the weak supervision as the guidance, the model’s VQA score increases from 62.31% to 62.83%. The results demonstrate that using weak supervision labels preserves potentially useful documents, aiding the Answerer in accurately answering questions.
Table 5. Ablation study on different methods of pseudo-labeling.
Model predictions Weak supervision labels Accuracy (%)
$\checkmark$ 62.31
$\checkmark$ $\checkmark$ 62.83
Effect of key knowledge documents ranges and quantities. In Tab. 6, we evaluate key knowledge document selection using various numbers of candidate documents and selected documents. From the results, we have the following findings: (1) As the number of selected documents increases, the model’s performance improves. This indicates that using more documents to train and test contributes to answering questions. (2) Using more documents for training can improve the performance a lot (the 2nd line $v.s.$ the last line). However, using more documents for testing has almost no improvement (the 3rd line $v.s.$ 4th line). (3) When the number of candidate documents increases, the model’s performance improves. The result demonstrates that low-ranked documents based on DPR scores may contain useful information for question answering. It is necessary for the model to select key knowledge documents.
Table 6. Ablation study on different numbers of candidate documents and selected documents.
$K_{candidate}$ $K_{train}$ $K_{test}$ Accuracy (%)
5 1 1 57.90
5 1 5 58.32
10 1 1 58.61
10 1 5 59.40
10 5 5 61.86
15 5 5 62.31
30 5 5 62.83
30 5 1 61.62
Effect of different vision information in our framework. In Tab. 7, we explore the impact of different vision information under different settings. We directly finetune BLIP2 without using knowledge and we find that the model using LLM-aligned query features outperforms the model using visual contexts, improving the accuracy from 41.82% to 55.11%. Using both can bring the performance improvement a little. In our framework, using query features improves the VQA score from 52.93% to 62.83%. These results demonstrate that introducing LLM-aligned visual features can help the model understand the visual details better, which contributes to improving the model’s performance. Using both degrades the performance a little.
Table 7. Ablation study on different vision information in our framework.
Models Query features Context Accuracy (%)
BLIP2 (fine-tuned) w/o knowledge $\checkmark$ 55.11
$\checkmark$ 41.82
$\checkmark$ $\checkmark$ 55.43
Boter (Ours) $\checkmark$ 62.83
$\checkmark$ 52.93
$\checkmark$ $\checkmark$ 62.52
Table 8. Ablation study on different documents selection in Answerer fine-tuning.
Knowledge Selection Accuracy (%)
Training Inference
DPR Selector 62.31
Selector DPR 60.75
Selector Selector 62.83
Figure 3. Qualitative results on the test split of OK-VQA. We compared our method with a model that fine-tunes BLIP2 with knowledge ranked by DPR. The middle segment of the graph represents knowledge from various methods used to answer questions. On the right side of the graph, different answers are depicted when using distinct knowledge. Green and red colors indicate whether the selected final answer is correct.
Effect of different knowledge documents selection in Answerer fine-tuning. Tab. 8 compares the Answerer fine-tuning with different document selection strategies. The results show that our framework performs optimally when utilizing Selector in both Answerer training and inference. This is likely because the Selector provides more informative key knowledge documents and using both Selector ensures the consistency between the training domain and testing domain.
4.4. Qualitative Analysis
In Fig. 3, We present a case study comparing our method with a model that fine-tunes BLIP2 using knowledge ranked by DPR. In the first case, top-ranked knowledge documents from DPR misguide the model, resulting in incorrect predictions. However, our method’s Selector chooses key knowledge documents that aid in predicting correct answers. In the second case, each knowledge document from DPR contains irrelevant information, leading to an incorrect final answer. Despite the top-1 document from the Selector resulting in a wrong answer, our method identifies other key knowledge documents for generating correct answers. Through majority voting, the final selected answer is correct. These cases demonstrate our method’s ability to extract informative knowledge from retrieved documents to support accurate question answering.
5. CONCLUSION
In this paper, we propose Boter, a novel framework to address the task of knowledge-based VQA. We adopt the multimodal language model to construct two modules: a Selector and an Answerer. The Selector selects key knowledge from the retrieved knowledge documents and the Answerer reasons on the key knowledge to answer questions. Through a cycle training method, we ensure consistent improvement of both modules’ abilities in knowledge selection and question answering. Compared with state-of-the-art methods, Boter achieves better performance on a challenging open-domain knowledge-based VQA benchmark (OK-VQA) and we conduct a comprehensive analysis to evaluate the effectiveness of our method.

### v3

@@
-Boter: Bootstrapping Knowledge Selection and Question Answering for Knowledge-based VQA
-Abstract.
-Knowledge-based Visual Question Answering (VQA) requires models to incorporate external knowledge to respond to questions about visual content.
-Previous methods mostly follow the ”retrieve and generate” paradigm.
-Initially, they utilize a pre-trained retriever to fetch relevant knowledge documents, subsequently employing them to generate answers.
-While these methods have demonstrated commendable performance in the task, they possess limitations: (1) they employ an independent retriever to acquire knowledge solely based on the similarity between the query and knowledge embeddings, without assessing whether the knowledge document is truly conducive to helping answer the question; (2) they convert the image into text and then conduct retrieval and answering in natural language space, which may not ensure comprehensive acquisition of all image information.
-To address these limitations, we propose Boter, a novel framework designed to bootstrap knowledge selection and question answering by leveraging the robust multimodal perception capabilities of the Multimodal Large Language Model (MLLM).
-The framework consists of two modules: Selector and Answerer, where both are initialized by the MLLM and parameter-efficiently finetuned in a simple cycle: find key knowledge in the retrieved knowledge documents using the Selector, and then use them to finetune the Answerer to predict answers; obtain the pseudo-labels of key knowledge documents based on the predictions of the Answerer and weak supervision labels, and then finetune the Selector to select key knowledge; repeat.
+Self-Bootstrapped Visual-Language Model for Knowledge Selection and Question Answering
+Abstract
+While large visual-language models (LVLM) have shown promising results on traditional visual question answering benchmarks, it is still challenging for them to answer complex VQA problems which requires diverse world knowledge.
+Motivated by the research of retrieval-augmented generation in the field of natural language processing, we use Dense Passage Retrieval (DPR) to retrieve related knowledge to help the model answer questions.
+However, DPR conduct retrieving in natural language space, which may not ensure comprehensive acquisition of image information.
+Thus, the retrieved knowledge is not truly conducive to helping answer the question, affecting the performance of the overall system.
+To address this issue, we propose a novel framework that leverages the visual-language model to select the key knowledge retrieved by DPR and answer questions.
+The framework consists of two modules: Selector and Answerer, where both are initialized by the LVLM and parameter-efficiently finetuned by self-bootstrapping: find key knowledge in the retrieved knowledge documents using the Selector, and then use them to finetune the Answerer to predict answers; obtain the pseudo-labels of key knowledge documents based on the predictions of the Answerer and weak supervision labels, and then finetune the Selector to select key knowledge; repeat.
 Our framework significantly enhances the performance of the baseline on the challenging open-domain Knowledge-based VQA benchmark, OK-VQA, achieving a state-of-the-art accuracy of 62.83%.
-We conduct comprehensive ablations of our framework, including Selector and Answerer ablations, the method of cycle training, varying key knowledge document numbers, the impact of vision information, and so on.
-Knowledge-based VQA, Selector and Answerer, Multimodal Large Language Model, Cycle training
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06
-Figure 1.
-Comparison with previous paradigms.
-(a) Previous KB-VQA paradigms apply captioning tools to transform the image into a visual context.
-They use a separate pre-trained knowledge retriever to retrieve knowledge and then use them to predict answers in natural language space.
-However, the top-ranked document contains noise information and misleads the model to answer a wrong answer ”surfing”.
-(b) The proposed paradigm aims to enhance the utilization of visual information to perform knowledge selection and question-answering consistency.
-We bootstrap the selector module and answerer module each other in a cycle style: where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector.
-The selector successfully finds key knowledge documents that are low-ranked but informative and the answerer predicts the right answer ”freestyle”.
-1. Introduction
-Visual Question Answering (VQA) (Antol et al., 2015) is a task that entails answering questions based on an image.
-Conventional VQA benchmarks (Goyal et al., 2017; Hudson and Manning, 2019) primarily focus on addressing straightforward questions that only necessitate visual perception and recognition.
-However, in real-world scenarios, many questions require broader world knowledge and common sense for accurate responses.
-Consequently, the Knowledge-based VQA (KB-VQA) benchmark is introduced to facilitate the development of models suitable for real-world applications.
-In contrast to previous VQA tasks, KB-VQA presents a more challenging setting where questions grounded in visual content demand answers by referencing external knowledge.
-Early studies (Wang et al., 2017a, b) offer a pre-defined knowledge base for all questions.
-More recently, the Outside-Knowledge Visual Question Answering (OK-VQA) (Marino et al., 2019) benchmark is introduced to emphasize the utilization of open-domain world knowledge to answer questions, rather than relying on a predefined knowledge base.
-This approach simulates a more realistic and human-like interaction with the visual world.
-To tackle the knowledge-based Visual Question Answering (VQA) task, a common approach involves retrieving relevant explicit knowledge and leveraging it to answer questions.
-Existing KB-VQA methods (Marino et al., 2021; Gardères et al., 2020; Lin and Byrne, 2022; Gui et al., 2021; Lin et al., 2022, 2023) often utilize the pre-trained retriever (e.g., DPR (Karpukhin et al., 2020)) to retrieve the related knowledge from diverse explicit knowledge bases, such as ConceptNet (Speer et al., 2017), Wikipedia (Vrandečić and Krötzsch, 2014), and Google Search Corpus (Luo et al., 2021).
-Subsequently, they utilize the retrieved knowledge to generate answers based on pre-trained language models.
-In addition to methods utilizing explicit knowledge, some approaches (Gui et al., 2021; Lin et al., 2022; Hu et al., 2022a; Shao et al., 2023) employ GPT-3 (Brown et al., 2020), a large-scale pre-trained language model, to generate implicit textual knowledge relevant to the question.
-With the explicit and implicit knowledge, the models have achieved good results.
-In addition to augmenting various knowledge bases, two critical components should be considered to obtain good performance: vision information mining and knowledge selection.
-To obtain vision information, many traditional methods transform the image into raw texts, such as captions (Hu et al., 2022a; Lin et al., 2022; Gui et al., 2021), tags (Yang et al., 2022), and OCR (Gao et al., 2022; Lin and Byrne, 2022).
-They use textual visual information in the process of retrieving and answering, which leads to the underutilization of visual information.
-On the other hand, to obtain external knowledge, many methods simply use a pre-trained retriever to obtain the relevant knowledge, without considering how to make the answer selector work better with the answer generator.
-The retrieved documents are only ranked based on the similarity between query and knowledge embeddings, which cannot guarantee that the selected documents can genuinely aid in producing correct answers.
-To address the aforementioned limitations, we introduce a novel framework, Boter, designed to select key knowledge documents and answer questions using the Multimodal Large Language Model (MLLM).
-We aim to enhance the utilization of visual information by leveraging the MLLM’s multimodal perception capabilities.
-Built upon the MLLM, our framework comprises two modules: a Selector and an Answerer.
+Our code is publicly available at https://github.com/haodongze/Self-KSel-QAns.
+Self-Bootstrapped Visual-Language Model for Knowledge Selection and Question Answering
+Dongze Hao1,2, Qunbo Wang1, Longteng Guo1, Jie Jiang1, Jing Liu1,2* 1Institute of Automation, Chinese Academy of Sciences, 2School of Artificial Intelligence, University of Chinese Academy of Sciences {haodongze2021,qunbo.wang}@ia.ac.cn, {longteng.guo,jie.jiang,jliu}@nlpr.ia.ac.cn thanks: Jing Liu is the corresponding author
+1 Introduction
+Recently, there has been an impressive advancement in large visual-language models (LVLM) Li et al.
+(2023); Alayrac et al.
+(2022); Liu et al.
+(2023); Dai et al.
+(2023).
+They usually use a mapping network to inject visual features into the semantic space of the large language model Brown et al.
+(2020); Zhang et al.
+(2022); Touvron et al.
+(2023); vic (2023); Touvron et al.
+(2023) and demonstrate strong capabilities on multimodal perception and reasoning.
+Thus, they achieve significant progress in conventional visual question answering benchmarks Antol et al.
+(2015); Goyal et al.
+(2017); Hudson and Manning (2019) which primarily focus on addressing straightforward questions that only necessitate visual perception and recognition.
+However, it is still challenging for the LVLMs to answer visual questions which require broader world knowledge and common sense Wang et al.
+(2017); Marino et al.
+(2019); Schwenk et al.
+(2022).
+Motivated by the research of retrieval-augmented generation Karpukhin et al.
+(2020a) in the field of natural language processing, we use Dense Passage Retrieval (DPR) to retrieve related world knowledge to help the model answer questions.
+However, when using DPR, we need to transform the image into texts to retrieve the related knowledge, which leads to the underutilization of visual information.
+Thus, the retrieved knowledge may be unfaithful and affects the model performance.
+To address the issue, we consider the LVLM as the knowledge selector to find helpful knowledge from candidate retrieved knowledge by DPR.
+Then the selected knowledge is fed into the LVLM to predict the answer.
+In this paper, we introduce a novel framework where we adopt the large visual-language model to perform knowledge selection and question answering.
+Our framework comprises two modules: a Selector and an Answerer.
 We train two modules by repeating the following process: the Selector first identifies important knowledge from the candidate knowledge documents retrieved by the pre-trained retriever; then, the Answerer takes the key knowledge documents as the input knowledge and is finetuned to generate the answer; next, we generate pseudo-labels of key knowledge documents according to the Answerer’s predictions and weak supervision labels; finally, we refine the Selector to assess the relevance of retrieved knowledge documents in answering the question.
-This strategy of cycle training bootstraps the ability of knowledge selection and answer generation consistently, enabling the model to accurately respond to knowledge-intensive questions.
-We conduct extensive experiments on the open-domain knowledge-based VQA benchmark (OK-VQA (Marino et al., 2019)) to validate the effectiveness of the proposed framework, where our method largely outperforms the baseline and achieves the state-of-the-art performance of 62.83%, only finetuning 0.16% parameters with LoRA (Hu et al., 2022b).
-We also conduct comprehensive ablations to validate the impact of different components of the proposed framework, including the Effect of Selector and Answerer, cycle training of the framework, varying the number of key knowledge documents, the impact of vision information, and so on.
+This strategy of self-bootstrapping enhances the ability of knowledge selection and answer generation consistently, enabling the model to accurately respond to knowledge-intensive questions.
+We conduct extensive experiments on the open-domain knowledge-based VQA benchmark (OK-VQA Marino et al.
+(2019)) to validate the effectiveness of the proposed framework, where our method largely outperforms the baseline and achieves the state-of-the-art performance of 62.83%, only finetuning 0.16% parameters with LoRA Hu et al.
+(2022a).
+We also conduct comprehensive ablations to validate the impact of different components of the proposed framework, including the Effect of Selector and Answerer, cycle training of the framework, varying the number of key knowledge documents, and so on.
 Our contributions are summarized as follows:
-  • We introduce a novel framework for consistently selecting key knowledge documents and answering questions using the multimodal language model.
-  • We propose a new cycle-training method to bootstrap Selector and Answerer each other, where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector.
-  • We achieve a state-of-the-art performance of 62.83% on the OK-VQA dataset, surpassing the previous state-of-the-art method. Notably, this improvement is achieved by fine-tuning only 0.16
-  • We present a comprehensive analysis to assess the impact of various components within the proposed framework.
-2. RELATED WORK
-Visual Question Answering (VQA).
-VQA presents a formidable challenge in visual-language understanding, requiring models to respond to visually-grounded questions.
-Conventional VQA benchmarks (Goyal et al., 2017; Hudson and Manning, 2019) primarily focus on basic visual perception and reasoning tasks.
-Numerous studies have achieved promising results on these benchmarks by incorporating: (1) improved visual features (Anderson et al., 2017; Jiang et al., 2020; Shen et al., 2021; Zhang et al., 2021); (2) attention-based cross-modality fusion techniques (Hu et al., 2017; Kim et al., 2018; Li et al., 2019; Yu et al., 2019); (3) the pretrain-finetune paradigm (Tan and Bansal, 2019; Lu et al., 2019; Zhang et al., 2021; Li et al., 2022, 2020; Wang et al., 2022).
-In addition to these general-purpose VQA tasks, the knowledge-based VQA task (Wang et al., 2017a; Marino et al., 2019) is introduced to advance models’ understanding of the open-ended world.
-This task leverages world knowledge to answer visual questions, thus enabling models to grasp the broader context of visual queries.
-Figure 2.
-Our Boter framework consists of two modules: a Selector and an Answerer.
+  • We introduce a novel framework that leverages the large visual-language model to select key knowledge and use them to answer questions, respectively.
+  • We propose a new self-bootstrap learning method to train the Selector and Answerer, where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector.
+  • We achieve a state-of-the-art performance of 62.83% on the OK-VQA dataset, surpassing the previous state-of-the-art method.
+Notably, this improvement is achieved by fine-tuning only 0.16% of parameters using LoRA.
+2 Related work
+Large Visual-Language Models.
+Recently, large visual-language models Li et al.
+(2023); Alayrac et al.
+(2022); Liu et al.
+(2023); Dai et al.
+(2023) have demonstrated remarkable visual-language understanding and reasoning capabilities, owing to the advancement of larger language models Brown et al.
+(2020); Zhang et al.
+(2022); Touvron et al.
+(2023); vic (2023); Touvron et al.
+(2023).
+These methods typically consist of a frozen visual encoder Radford et al.
+(2021), a visual-language connector Li et al.
+(2023), and a large language model Chung et al.
+(2022); Zhang et al.
+(2022); vic (2023).
+The models are firstly pre-trained on large-scale visual-text datasets to align visual features to the language embedding space.
+After pretraining, the large language model can understand the visual details.
+Then, the model is finetuned to adapt to various visual-language tasks.
+In this study, we adopt BLIP2, one of the widely used models, as our backbone for bootstrapping knowledge selection and question answering with it.
+Figure 1: Our framework consists of two modules: a Selector and an Answerer.
 Selector (left) selects the top-T knowledge documents for the Answerer (right), and the Answerer focuses on important knowledge information to predict answers.
@@
 Knowledge-based VQA.
-Early benchmarks (Wang et al., 2017a, 2015) provide fixed knowledge bases for the model to retrieve and answer questions.
-Recently, the open-domain knowledge-based VQA benchmark (OK-VQA (Marino et al., 2019)) is proposed, which provides no extra knowledge bases and the model needs to find proper knowledge bases to help answer questions.
-Recent studies (Gardères et al., 2020; Wu et al., 2022; Lin and Byrne, 2022; Gui et al., 2021) have explored various open-domain world knowledge sources, such as ConceptNet (Speer et al., 2017), Wikipedia (Vrandečić and Krötzsch, 2014), Google Search Corpus (Luo et al., 2021).
+Conventional VQA benchmarks Goyal et al.
+(2017); Hudson and Manning (2019) primarily focus on basic visual perception and reasoning tasks and numerous studies have achieved promising results on these benchmarks Anderson et al.
+(2017); Zhang et al.
+(2021); Tan and Bansal (2019); Lu et al.
+(2019); Li et al.
+(2022); Wang et al.
+(2022).
+Different from them, the knowledge-based VQA task Wang et al.
+(2017); Marino et al.
+(2019); Schwenk et al.
+(2022) requires models to incorporate diverse world knowledge to respond to questions about visual content, which is more challenging.
+Recent studies Gardères et al.
+(2020); Wu et al.
+(2022); Lin and Byrne (2022); Gui et al.
+(2021) have explored various open-domain world knowledge sources, such as ConceptNet Speer et al.
+(2017), Wikipedia Vrandečić and Krötzsch (2014), Google Search Corpus Luo et al.
+(2021).
 They retrieve the relevant knowledge documents from the knowledge bases and integrate them into the answering model to generate predictions.
-Except for using explicit knowledge, some methods also take GPT-3 (Brown et al., 2020) as an implicit knowledge producer.
-They either prompt GPT-3 with in-context examples to predict answers directly (Yang et al., 2022; Hu et al., 2022a; Shao et al., 2023), or use GPT-3 to generate answer candidates with evidence serving as textual implicit knowledge bases (Gui et al., 2021; Lin et al., 2022), leading to significant performance improvements.
-Different from previous approaches that primarily enhance performance by incorporating more knowledge sources, our focus lies in effectively leveraging visual information and retrieved knowledge.
-We employ a multimodal large language model to enhance visual understanding and construct a Selector and an Answerer to consistently select key knowledge and answer questions.
-Multimodal large language model.
-Recently, multimodal large language models (Li et al., 2023; Alayrac et al., 2022; Liu et al., 2023; Dai et al., 2023) have demonstrated remarkable visual-language understanding and reasoning capabilities, owing to the advancement of larger language models (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023; vic, 2023; Touvron et al., 2023).
-These methods typically consist of a frozen visual encoder (Radford et al., 2021), a visual-language connector (Li et al., 2023), and a large language model (Chung et al., 2022; Zhang et al., 2022; vic, 2023).
-The models are firstly pre-trained on large-scale visual-text datasets to align visual features to the language embedding space.
-After pretraining, the large language model can understand the visual details.
-Then, the model is finetuned to adapt to various visual-language tasks.
-In this study, we adopt BLIP2, one of the widely used models, as our backbone for understanding visual details and bootstrap knowledge selection and question answering with it.
-3. Method
-In this section, we first introduce the preliminaries of Knowledge Retrieval and MLLM, which are the foundation of our framework.
+Except for using explicit knowledge, some methods also take GPT-3 Brown et al.
+(2020) as an implicit knowledge producer.
+They either prompt GPT-3 with in-context examples to predict answers directly Yang et al.
+(2022); Hu et al.
+(2022b); Shao et al.
+(2023), or use GPT-3 to generate answer candidates with evidence serving as textual implicit knowledge bases Gui et al.
+(2021); Lin et al.
+(2022), leading to significant performance improvements.
+Different from these approaches, we employ a large visual-language model to select key retrieved knowledge and reason on the knowledge to answer questions.
+3 Method
+In this section, we first introduce the preliminaries of Knowledge Retrieval and LVLM, which are the foundation of our framework.
 Then, we present the design of the Selector and Answerer for knowledge selection and question answering on knowledge respectively.
-Finally, we illustrate the cycle training method of two designed modules.
-3.1. Preliminaries
+Finally, we illustrate the self-bootstrap training method of two designed modules.
+3.1 Preliminaries
 Knowledge Retrieval.
-Following previous knowledge-based VQA methods (Gao et al., 2022; Lin and Byrne, 2022; Si et al., 2023), we adopt the dense passage retrieval (DPR) (Karpukhin et al., 2020) to retrieve the knowledge documents.
-We transform the visual image $I_{i}$ into texts through the following steps:
-  • The pretrained captioning model Oscar+ (Zhang et al., 2021) is adopted to obtain the image caption $C_{i}$.
-  • The pretrained detection model VinVL (Zhang et al., 2021) is adopted to extract the text description of objects and attributes $L_{i}$.
-  • The Google OCR (Optical Character Recognition) APIs are adopted to extract the text strings $\{O_{i}\}$ of the image.
-Consequently, the visual contexts of the image $I$ can be represented as $V_{i}=(C_{i},L_{i},O_{i})$.
-We combine the question $Q_{i}$ and the visual contexts $V_{i}$ as a query $q_{i}=(Q_{i},V_{i})$ to retrieve the relevant documents from the knowledge base $D$.
-DPR encodes the query $q_{i}$ and all the knowledge documents $\mathcal{D}$ into embeddings $\mathbf{q}_{i}$ and $\{\mathbf{d}_{1},\mathbf{d}_{2},...,\mathbf{d}_{n}\}$ with different BERT-like (Devlin et al., 2018) encoders as follows:
-(1) $\mathbf{q}_{i}=BERT_{q}(q_{i}),\mathbf{d}_{i}=BERT_{d}(D_{i}),$
-We compute the similarity scores $sim(q_{i},D_{j})=\mathbf{q}_{i}^{T}\cdot\mathbf{d}_{j}$ and exploit FAISS (Johnson et al., 2019) to index Top-k related knowledge documents $\mathcal{P}_{i}=\{P_{i,1},P_{i,2},...,P_{i,k}\}$ for $i$-th query.
-Multimodal Large Language Model.
-Recently, multimodal large language models (Li et al., 2023; Alayrac et al., 2022; Liu et al., 2023) have demonstrated significant advancements in visual-language understanding.
-The architecture of these models typically comprises a frozen image encoder (Dosovitskiy et al., 2020; Fang et al., 2023), a visual-language connector (Li et al., 2023), and a pre-trained language model (Chung et al., 2022; Zhang et al., 2022; vic, 2023).
+We adopt the Dense Passage Retrieval (DPR) Karpukhin et al.
+(2020b) to retrieve the knowledge documents.
+We transform the image into raw texts composed of captions, objects, attributes, and OCR (Optical Character Recognition).
+Then we compute the similarity scores between the query and knowledge documents $sim(q_{i},D_{j})=\mathbf{q}_{i}^{T}\cdot\mathbf{d}_{j}$ and exploit FAISS Johnson et al.
+(2019) to index Top-k related knowledge documents $\mathcal{P}_{i}=\{P_{i,1},P_{i,2},...,P_{i,k}\}$ for $i$-th query.
+Large Visual-Language Model.
+In our work, both knowledge selection and question-answering modules adopt BLIP-2 Li et al.
+(2023) as the backbone.
+The architecture of BLIP-2 comprises a frozen image encoder Dosovitskiy et al.
+(2020); Fang et al.
+(2023), a Q-Former Li et al.
+(2023), and a pre-trained language model Chung et al.
+(2022).
 Given an image $I_{i}$, the frozen image encoder outputs a set of visual features $\{\mathbf{h}_{i,1},\mathbf{h}_{i,2},...,\mathbf{h}_{i,m}\}$.
-Visual-language connector takes extracted visual features as input, and outputs language-aligned visual features $\{\mathbf{v}_{i,1},\mathbf{v}_{i,2},...,\mathbf{v}_{i,l}\}$.
+Q-Former takes extracted visual features as input, and outputs language-aligned visual features $\{\mathbf{v}_{i,1},\mathbf{v}_{i,2},...,\mathbf{v}_{i,l}\}$.
 These visual features are concatenated with the textual word embeddings, which are fed into the language model for generation.
-Through pre-training on large-scale image-caption datasets, the visual-language connector can effectively project visual features into the feature space of the Language Large Model (LLM).
-In our work, both knowledge selection and question-answering modules adopt BLIP-2 (Li et al., 2023) as the backbone.
+Through pre-training on large-scale image-caption datasets, Q-Former can effectively project visual features into the feature space of the Language Large Model (LLM).
 We freeze the visual encoder and Q-former during training.
-We train the fully connected layer and use LoRA (Hu et al., 2022b) to finetune the LLM (only finetune 0.16% of total parameters).
-3.2. Selector and Answerer
-For the knowledge-based VQA task, previous methods (Lin and Byrne, 2022; Gui et al., 2021; Lin et al., 2022) separately use DPR to retrieve relevant knowledge documents from the knowledge base, and then directly use them to help produce the answer for the given question.
-While these methods have demonstrated the effectiveness of leveraging external knowledge, the separate retriever is not inherently aligned with the answerer.
-The retrieved knowledge documents are ranked based on similarity scores between the query and knowledge embeddings, which may not always be useful for answering a specific question.
-Additionally, these methods transform the image into textual information and conduct retrieval and answering in the language space, resulting in suboptimal utilization of visual information.
-We adopt a multimodal large language model, BLIP-2 (Li et al., 2023), to enhance the utilization of visual information.
-We assign BLIP-2 two roles: a Selector and an Answerer.
-The Selector is responsible for selecting key knowledge from the retrieved documents that truly assist in answering questions, while the Answerer predicts the answer based on the selected knowledge.
-This integrated approach ensures better alignment between knowledge selection and question answering and enhances the model’s ability to answer questions.
+We train the fully connected layer and use LoRA Hu et al.
+(2022a) to finetune the LLM (only finetune 0.16% of total parameters).
+3.2 Selector and Answerer
 Selector.
-After obtaining the Top-k knowledge documents using DPR for the $i$-th sample, we aim to choose $t$ most important knowledge documents from the retrieved documents.
-where $t$ is smaller than $k$.
+After obtaining the Top-k knowledge documents using DPR for the $i$-th sample, we aim to choose $t$ most important knowledge documents from the retrieved documents, where $t$ is smaller than $k$.
 As shown in Fig.
-2, we firstly use the frozen image encoder and Q-former to extract the image features $\mathbf{V}_{i}$, where these features are extracted once and then used by the Selector and the Answerer.
+1, we first use the frozen image encoder and Q-former to extract the image features $\mathbf{V}_{i}$, where these features are extracted once and then used by the Selector and the Answerer.
 Then image features $\mathbf{V}_{i}$ are fed into the independent fully-connected layer to obtain the visual embeddings $\mathbf{E}^{v}_{i}$.
-We concatenate the question, a retrieved knowledge document, and the Selection prompt ”Does the retrieved knowledge document provide the key information to help answer the question?” into one sentence $S$.
-Next, visual embeddings $\mathbf{E}^{v}_{i}$ and the text are concatenated and fed into the LLM (Flan-T5 (Chung et al., 2022) is adopted in our work).
+We concatenate the question, a retrieved knowledge document, and the Selection prompt "Does the retrieved knowledge document provide the key information to help answer the question?" into one sentence $S$.
+Next, visual embeddings $\mathbf{E}^{v}_{i}$ and the text are concatenated and fed into the LLM (Flan-T5 Chung et al.
+(2022) is adopted in our work).
 Last, we use the probability of generating the word ‘yes’ as the score of each retrieved knowledge document $P_{i,j}$, denoted as $s_{i,j}=\bm{LLM}(concat(\mathbf{E}^{v}_{i},S_{i}))$, and we select top-t documents $\hat{\mathcal{P}}_{i}=\{\hat{P}_{i,1},\hat{P}_{i,2},...,\hat{P}_{i,t}\}$ based on the scores.
 The Selector can be conceptualized as follows:
-(2) $\hat{\mathcal{P}}_{i}==Selector(I_{i},Q_{i},\mathcal{P}_{i}),|\hat{\mathcal{P}}_{i}|=t$
+$\hat{\mathcal{P}}_{i}==Selector(I_{i},Q_{i},\mathcal{P}_{i}),|\hat{\mathcal{P}}_{i}|=t$ (1)
 Answerer.
@@
 As shown in Fig.
-2, we process the same image features to obtain the different visual embeddings $\mathbf{E}^{v}_{i}$ via the fully-connected layer of the Answerer.
-Next, we concatenate the question and the knowledge into one sentence $S^{\prime}$ using the template ”Question: {} Knowledge: {} Answer: ”.
+1, we process the same image features to obtain the different visual embeddings $\mathbf{E}^{v}_{i}$ via the fully-connected layer of the Answerer.
+Next, we concatenate the question and the knowledge into one sentence $S^{\prime}$ using the template "Question: {} Knowledge: {} Answer: ".
 We concatenate the visual embeddings and the text, which are fed into the LLM with different LoRA parameters to get the answer.
@@
 The Answerer can be conceptualized as follows:
-(3) $a_{i}=Answerer(I_{i},Q_{i},\hat{\mathcal{P}}_{i})$
+$a_{i}=Answerer(I_{i},Q_{i},\hat{\mathcal{P}}_{i})$ (2)
 Then the final answer is based on the majority vote. We also tried different knowledge reasoning methods, such as concatenating (the results can be seen in the ablation study).
@@
   end for
-3.3. Cycle Training
+3.3 Self-Bootstrap Learning
 To enable the Selector and Answerer to select key knowledge and answer questions, we bootstrap them with each other in a style of cycle training.
@@
 Then we finetune the Answerer with LoRA under the supervision of the ground truth answer $\mathcal{A}_{i}$:
-(4) $\mathbf{E}^{v}_{ans,i}=FC_{ans}(\mathbf{V}_{i}),$
+$\mathbf{E}^{v}_{ans,i}=FC_{ans}(\mathbf{V}_{i}),$ (3)
 $L_{ans}=-\sum_{j=1}^{t}\log LLM_{ans}(a_{i}^{*}|\mathbf{E}^{v}_{ans,i},Q_{i},\hat{P}_{i}^{j}),$
@@
 2 to predict answers based on each retrieved knowledge document $P_{i,j}$.
-Then we assign pseudo labels to the retrieved documents according to model predictions and weak supervision labels (Luo et al., 2021; Lin and Byrne, 2022; Lin et al., 2023).
-We use ”yes” and ”no” as pseudo labels, where label a document as positive knowledge if Answerer can output the correct answer using that document and the document contains any of the answers in $\mathcal{A}_{i}$.
-(5) $y_{i,j}=\begin{cases}\text{yes},\quad\textit{if}\ \ a_{i}=a_{i}^{*}\wedge P_{i,j}\ \ \text{contains an answer in}\ \ \mathcal{A}_{i}\\ \text{no},\quad else\end{cases}$
-Table 1.
-Performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset.
+Then we assign pseudo labels to the retrieved documents according to model predictions and weak supervision labels Luo et al.
+(2021); Lin and Byrne (2022); Lin et al.
+(2023).
+We use "yes" and "no" as pseudo labels, where label a document as positive knowledge if Answerer can output the correct answer using that document and the document contains any of the answers in $\mathcal{A}_{i}$.
+$y_{i,j}=\begin{cases}\text{yes},\quad\textit{if}\ \ a_{i}=a_{i}^{*}\wedge\\ \quad\quad P_{i,j}\ \ \text{contains an answer in}\ \ \mathcal{A}_{i}\\ \text{no},\quad else\end{cases}$ (4)
+Table 1: Performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset.
 Knowledge Sources: ConceptNet (C); Wikipedia (W); Google Search (GS); Google Images (GI).
 The best result in the table is bolded.
-Compared with the previous SOTA method Prophet (Shao et al., 2023) using GPT-3, our method achieves 62.83% performance, with a performance improvement of 1.73%.
-The results demonstrate the effectiveness of our method.
-Models Large Models $K_{train}$ $K_{test}$ Knowledge Resource Accuracy (%)
-BAN+AN (Marino et al., 2019) - - - W 25.6
-ConceptBERT (Gardères et al., 2020) - - - C 33.7
-KRISP (Marino et al., 2021) - - - C+W 38.4
-Visual Retriever-Reader (Luo et al., 2021) - 100 100 GS 39.2
-MAVEx (Wu et al., 2022) - - - W+C + GI 39.4
-TRiG(Ensemble) (Gao et al., 2022) T5-large (770M) 100 100 W 50.5
-KAT(Single) (Gui et al., 2021) T5-large (770M) 40 40 W + GPT-3 (175B) 53.1
-KAT(Ensemble) (Gui et al., 2021) T5-large (770M) 40 40 W + GPT-3 (175B) 54.4
-RA-VQA (Lin and Byrne, 2022) T5-large (770M) 5 50 GS 54.5
-REVIVE(Single) (Lin et al., 2022) T5-large (770M) 40 40 W+GPT-3 (175B) 56.6
-REVIVE(Ensemble) (Lin et al., 2022) T5-large (770M) 40 40 W+GPT-3 (175B) 58.0
-BLIP2 T5-XXL (Dai et al., 2023) BLIP2 T5-XXL (11B) - - BLIP2 54.7
-PICa (Yang et al., 2022) GPT-3 (175B) - - GPT-3 (175B) 48.0
-Flamingo (Alayrac et al., 2022) Flamingo (80B) - - Flamingo (80B) 57.8
-PromptCap (Hu et al., 2022a) GPT-3 (175B) - - GPT-3 (175B) 60.4
-Prophet (Shao et al., 2023) GPT-3 (175B) - - GPT-3 (175B)+MCAN 61.1
-Boter(Ours) BLIP2 T5-XL (3B) 5 5 GS 62.83
+The results show that our method achieves the state-of-the-art performance.
+Models Large Models $K_{train}$ $K_{test}$ Knowledge Resource Acc (%)
+BAN+AN Marino et al. (2019) - - - W 25.6
+ConceptBERT Gardères et al. (2020) - - - C 33.7
+KRISP Marino et al. (2021) - - - C+W 38.4
+Visual Retriever-Reader Luo et al. (2021) - 100 100 GS 39.2
+MAVEx Wu et al. (2022) - - - W+C + GI 39.4
+PICa Yang et al. (2022) GPT-3 (175B) - - GPT-3 48.0
+TRiG(Ensemble) Gao et al. (2022) T5-large (770M) 100 100 W 50.5
+KAT(Single) Gui et al. (2021) T5-large (770M) 40 40 W + GPT-3 53.1
+KAT(Ensemble) Gui et al. (2021) T5-large (770M) 40 40 W + GPT-3 54.4
+RA-VQA Lin and Byrne (2022) T5-large (770M) 5 50 GS 54.5
+REVIVE(Single) Lin et al. (2022) T5-large (770M) 40 40 W+GPT-3 56.6
+REVIVE(Ensemble) Lin et al. (2022) T5-large (770M) 40 40 W+GPT-3 58.0
+PromptCap Hu et al. (2022b) GPT-3 (175B) - - GPT-3 60.4
+Prophet Shao et al. (2023) GPT-3 (175B) - - GPT-3+MCAN 61.1
+FillingGap Wang et al. (2023) GPT-3 (175B) - - GPT-3 61.3
+SimpleBaseline Xenos et al. (2023) LLaMA 2 (13B) - - LLaMA 2 61.2
+Cola-FT Chen et al. (2024) FLAN-T5(11B) - - BLIP+OFA 62.4
+Flamingo Alayrac et al. (2022) Flamingo (80B) - - Pretrain 57.8
+InstructBLIP Dai et al. (2023) InstructBLIP Vicuna (7B) - - Pretrain 62.1
+Qwen-VL Bai et al. (2023) Qwen-VL(Qwen-7B) - - Pretrain 58.6
+MM-Reasoner Khademi et al. (2023) Flamingo (80B) - - GPT-4 60.8
+BLIP2 (fine-tuned) Li et al. (2023) BLIP2 T5-XL (3B) - - Pretrain 55.4
+RA-VQA-v2 Lin et al. (2023) BLIP2 T5-XL (3B) 5 5 GS 62.1
+PreFLMR Lin et al. (2024) BLIP2 T5-XL (3B) 5 5 GS 61.8
+Ours BLIP2 T5-XL (3B) 5 5 GS 62.8
 After obtaining the pseudo label of each retrieved knowledge document, we use the trainable $FC_{sel}$ layer to output the visual embeddings $\mathbf{E}^{v}_{sel,i}$.
@@
 Then we finetune the Selector with LoRA under the supervision of pseudo labels:
-(6) $\mathbf{E}^{v}_{sel,i}=FC_{sel}(\mathbf{V}_{i}),$
+$\mathbf{E}^{v}_{sel,i}=FC_{sel}(\mathbf{V}_{i}),$ (5)
 $L_{sel}=-\sum_{j=1}^{k}\log LLM_{sel}(y_{i,j}|\mathbf{E}^{v}_{sel,i},Q_{i},P_{i}^{j})$
@@
 During the inference stage, we utilize the Selector to choose key knowledge, and then instruct the Answerer to respond to questions based on this knowledge.
-4. EXPERIMENTS
-4.1. Experimental Setup
+4 Experiments
+4.1 Experimental Setup
 Dataset.
-We conducted extensive experiments on OK-VQA (Marino et al., 2019) to evaluate the effectiveness of our method.
+We conduct extensive experiments on OK-VQA Marino et al.
+(2019) to evaluate the effectiveness of our method.
 OK-VQA is a challenging open-domain knowledge-based VQA dataset that requires models to leverage various external knowledge sources to answer questions.
@@
 Due to no knowledge base being provided for OK-VQA, we need to choose the proper knowledge base for the dataset.
-In this paper, we adopt Google Search Corpus (Luo et al., 2021) as the knowledge base which is collected in the websites using the Google Search API.
-The dataset has been proven to be effective by previous methods (Luo et al., 2021; Lin and Byrne, 2022; Lin et al., 2023).
-We do not evaluate our method on A-OKVQA (Schwenk et al., 2022) like other retrieved-based methods (Gui et al., 2021; Lin et al., 2022; Lin and Byrne, 2022; Lin et al., 2023), as it emphasizes visually-grounded reasoning rather than knowledge retrieval and selection.
+In this paper, we adopt Google Search Corpus Luo et al.
+(2021) as the knowledge base which is collected in the websites using the Google Search API.
 Evaluation Metric.
-We use the standard VQA metric (Antol et al., 2015) to evaluate the performance of the model.
+We use the standard VQA metric Antol et al.
+(2015) to evaluate the performance of the model.
 Given the prediction of the question $a$ and the groudtruth answer set $\mathcal{A}$, the VQA accuracy is calculated as:
-(7) $Accuracy(a,\mathcal{A})=\min(\frac{\#A(a)}{3},1),$
+$Accuracy(a,\mathcal{A})=\min(\frac{\#A(a)}{3},1),$ (6)
 where the groudtruth answer set $\mathcal{A}$ is annotated by different humans, #$\#A(a)$ denotes the occurrence of $a$ in $\mathcal{A}$.
 Implementation Details.
-In our experiment, we adopt BLIP2 T5-XL (3B) (Li et al., 2023) to initialize the Selector and Retriever.
-We freeze the image encoder and Q-former, with both the Selector and Retriever sharing the same visual module.
-We finetune the fully connected layer and use LoRA (Hu et al., 2022b) to train the LLM.
+In our experiment, we adopt BLIP2 T5-XL (3B) Li et al.
+(2023) to initialize the Selector and Answerer.
+We freeze the image encoder and Q-former, with both the Selector and Answerer sharing the same visual module.
+We finetune the fully connected layer and use LoRA Hu et al.
+(2022a) to train the LLM.
 We use the default huggingface-PEFT setting: r=8, lora_alpha=32, lora_dropout=0.1.
@@
 We use the warm-up strategy which trains the model with an initial learning rate of 1e-4 and warm-up factor of 0.05 for 1000 steps and then utilizes a cosine annealing learning strategy with an initial learning rate of 1e-4 and a final learning rate of 0 after 10 epochs.
-We use top-30 knowledge documents retrieved by a pre-trained DPR (Karpukhin et al., 2020) as candidates for Selector and use the selected top-5 documents from the 30 documents for the Answerer to train and infer, denoted as $K_{candidate}=30,K_{train}=5,K_{test}=5$.
+We use top-30 knowledge documents retrieved by a pre-trained DPR Lin and Byrne (2022) as candidates for Selector and use the selected top-5 documents from the 30 documents for the Answerer to train and infer, denoted as $K_{candidate}=30,K_{train}=5,K_{test}=5$.
 We use 2 Nvidia A800 GPUs (80G) for all experiments.
-And our codes will be released upon paper acceptance.
-4.2. Overall Results
-We demonstrate the performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset in the Tab.
-1.
-From the results, we have the following observations: (1) Early models (BAN+AN (Marino et al., 2019), ConceptBERT (Gardères et al., 2020), KRISP (Marino et al., 2021), Visual Retriever-Reader (Luo et al., 2021), and MAVEx (Wu et al., 2022)) adopt the small visual-language model as backbones and inject various knowledge bases (ConceptNet, Wikipedia, Google Web Search and Google Images) into the model, achieving a VQA accuracy from 25.6% to 39.4%.
-Although these models use external knowledge to help produce answers, they have relatively low performance because of the weak knowledge reasoning ability of traditional backbones.
-(2) Recent methods (TRiG(Gao et al., 2022), KAT (Gui et al., 2021), RA-VQA (Lin and Byrne, 2022), and REVIVE (Lin et al., 2022)) introduce the large language model to improve the capacity for common-sense reasoning.
-They exploit T5-large (Chung et al., 2022) to reason on the knowledge for question answering.
-Compared with early methods, they have a higher accuracy ranging from 50.5% to 58.0%.
-(3) Some recent methods take a very large language model GPT-3 (Brown et al., 2020) as the implicit knowledge base for OK-VQA.
-PICa (Yang et al., 2022) is the first method to use GPT-3.
-It uses the style of in-context learning to prompt GPT-3 to answer questions directly.
-Inspired by PICa, some methods (Hu et al., 2022a; Shao et al., 2023) improve the performance of GPT-3 by generating question-guided captions and providing answer candidates, achieving higher VQA score on OK-VQA, about 61%.
-Meanwhile, some retrieved-based methods (Gui et al., 2021; Lin et al., 2022) use GPT-3 to generate implicit knowledge and combine them with the retrieved explicit knowledge, which improves the performance.
-(4) Recent large multimodal large language models have achieved high performance on OK-VQA without using external knowledge (e.g., Flamingo (Alayrac et al., 2022) (80B) has a zero-shot VQA score of 57.8%, BLIP2 T5-XXL (Dai et al., 2023) has a finetuning VQA score of 54.7%).
-(5) Our method aims to select key knowledge and answer questions utilizing the multimodal perception and reasoning ability of the multimodal large language model.
-By cycle training, we bootstrap Selector and Answerer to perform knowledge selection and question answering consistently.
-With a small number of selected knowledge documents, our method outperforms other state-of-the-art models, achieving an accuracy of 62.83%.
+4.2 Comparison with State-of-the-art Methods
+As shown in Tab.
+1, we can see that early models (BAN+AN Marino et al.
+(2019), ConceptBERT Gardères et al.
+(2020), KRISP Marino et al.
+(2021), Visual Retriever-Reader Luo et al.
+(2021), and MAVEx Wu et al.
+(2022)) have a weak performance, achieving a VQA accuracy from 25.6% to 39.4%.
+Recently, by introducing larger models (T5-large, GPT-3, LLaMA, Vicuna) and diverse knowledge resources (ConceptNet, Wikipedia, Google Web Search and Google Images), the performance has a significant performance improvement, achieving a VQA accuracy of 62.4%.
+Our method aims to augment the reasoning ability to answer knowledge-intensive questions of the large visual-language model.
+When directly finetuning BLIP2 T5-XL on OKVQA, the model has a low performance of 55.44%.
+By introducing external knowledge, the performance has a significant performance improvement.
+Different from RA-VQA-v2 Lin et al.
+(2023) and PreFLMR Lin et al.
+(2024), we do not train a multimodal retriever from scratch which requires expensive annotations and high computational costs.
+We directly leverage the large visual-language model to select key knowledge from the retrieved knowledge by DPR like the process of re-ranking.
+With the same knowledge resources (i.e., Google Search), our method achieves 62.8% accuracy, outperforming other state-of-the-art models.
 It is worth noting that we do not use GPT-3 and we only train the 0.16% parameters of the model.
-4.3. Ablation Study
+These results demonstrate the effectiveness of the proposed approach.
+4.3 Ablation Study
 We conduct the ablation studies to evaluate different components of our framework on OK-VQA.
@@
 2.
-From the results, we can observe: (1) Our framework, leveraging key knowledge documents selected by the Selector, consistently outperforms the Answerer when using the same number of documents retrieved by DPR.
+From the results, we can observe: our framework, leveraging key knowledge documents selected by the Selector, consistently outperforms the Answerer when using the same number of documents retrieved by DPR.
 We improve the performance by 2.14% and 1.88% with 1 and 5 test knowledge documents, compared to DPR-based retrieval.
@@
 These results demonstrate that top-ranked knowledge documents based on DPR scores are not optimal for question answering and our key knowledge selection module can identify relevant documents for accurate question answering, ensuring the coherence of knowledge retrieval and question-answering processes.
-(2) With an increase in the number of test knowledge documents to 10, the performance of the model DPR-based retrieval almost has no improvement and our method’s performance decreases.
-This suggests that additional documents may introduce more irrelevant knowledge, adversely affecting question-answering accuracy.
-Table 2.
-Ablation study on the Selector.
+Table 2: Comparison of our selector with different knowledge selection strategies.
 We select 5 knowledge documents from top-30 knowledge candidates retrieved by DPR.
-’DPR Score’ refers to selecting top-5 knowledge based on similarity scores.
-’Random Selection’ means randomly selecting 5 knowledge documents from 30 candidate knowledge documents.
-’Selector’ denotes choosing 5 key knowledge documents by the Selector.
-$K_{train}$ $K_{test}$ Knowledge Selection Accuracy (%)
+DPR Score refers to selecting top-5 knowledge based on similarity scores.
+Random Selection means randomly selecting 5 knowledge documents from 30 candidate knowledge documents.
+Selector denotes choosing 5 key knowledge documents by the Selector.
+$K_{train}$ $K_{test}$ Knowledge Selection Acc (%)
 5 1 Random Selection 50.45
@@
 5 5 Selector 62.83
-5 10 Random Selection 52.11
-5 10 DPR Score 60.93
-5 10 Selector 62.17
-Effect of Answerer.
+Effect of different knowledge reasoning methods of Answerer.
 In Tab.
-3, we present a comparison of our Answerer using different knowledge reasoning methods.
+3, we present a comparison of Answerer using different knowledge reasoning methods.
 The results show that the performance using the strategy of voting surpasses that of concatenating under different knowledge selection settings.
-We argue that directly combining all the knowledge documents into a lengthened document makes it difficult for the Answerer to reason on them, which is easily influenced by noisy information.
-In contrast, it is easier for the Answerer to reason on each document to predict the answer.
+We argue that directly combining all the knowledge documents into a lengthened document makes it difficult for Answerer to reason on them, which is easily influenced by noisy information.
+In contrast, it is easier for Answerer to reason on each document to predict the answer.
 Simple voting can choose the best answer.
-Table 3.
-Ablation study on different knowledge reasoning methods of Answerer.
-’Concatenating’ denotes that we combine the key knowledge documents into one sentence and feed it into the Answerer to predict the final answer.
-’Voting’ means that we feed different key knowledge documents into the Answerer to predict different answers and choose the best answer based on majority voting.
-Method Backbone Accuracy (%)
-Concatenating BLIP2 (fine-tuned) w knowledge from DPR 59.11
+Table 3: Effect of different knowledge reasoning methods of Answerer.
+Concatenating denotes that we combine the key knowledge documents into one sentence and feed it into Answerer to predict the final answer.
+Voting means that we feed different key knowledge documents into Answerer to predict different answers and choose the best answer based on majority voting.
+Method VQA Model Acc (%)
+Concatenating BLIP2 (fine-tuned) 59.11
 Voting 60.69
-Concatenating Boter (Ours) w knowledge from Selector 62.06
+Concatenating Ours 62.06
 Voting 62.83
-Effect of cycle training.
-To evaluate the effectiveness of our cycle-training method, we compare the method with the strategy of independent training of two modules.
+Effect of the self-bootstrap learning method.
+To evaluate the effectiveness of our self-bootstrap learning method, we compare the method with the strategy of independent training of two modules.
 We finetune the Answerer with the knowledge documents retrieved by DPR as the baseline.
-Independent training means that we train the Selector and the Answerer respectively.
-Initially, we train the Answerer module utilizing knowledge documents retrieved by DPR.
-Subsequently, employing the trained Answerer, we generate answers for each retrieved knowledge document, thereby generating pseudo-labels for the retrieved knowledge.
-We then proceed to train the Selector module supervised by these pseudo-labels.
-Finally, we conduct finetuning of the Answerer once more, incorporating new knowledge documents selected by the Selector.
+Independent training means that we take two passes in answerer training and one pass for selector training on the entire dataset.
+Cycle training means that we train the answerer and selector on each batch data of the dataset simultaneously.
 The results in Tab.
-4 show that the model with cycle training outperforms the model with iterative refinement by 3.81%.
+4 show that the model with cycle training outperforms the model with independent training by 3.81%.
 The VQA score of using independent training is even lower than the baseline.
 These results demonstrate that our cycle training method can effectively boost the Selector and Answerer each other, which makes the model find key knowledge documents and leverage the knowledge to answer questions.
-Table 4. Ablation study on different training methods of our framework.
-Methods Accuracy (%)
+Table 4: Effect of the self-bootstrap learning method.
+Method Acc (%)
 Baseline 60.69
@@
 The results demonstrate that using weak supervision labels preserves potentially useful documents, aiding the Answerer in accurately answering questions.
-Table 5. Ablation study on different methods of pseudo-labeling.
-Model predictions Weak supervision labels Accuracy (%)
+Table 5: Ablation study on different methods of pseudo-labeling.
+Model predictions Weak supervision labels Acc (%)
 $\checkmark$ 62.31
 $\checkmark$ $\checkmark$ 62.83
-Effect of key knowledge documents ranges and quantities.
+Effect of key knowledge documents selection ranges and quantities.
 In Tab.
@@
 It is necessary for the model to select key knowledge documents.
-Table 6. Ablation study on different numbers of candidate documents and selected documents.
-$K_{candidate}$ $K_{train}$ $K_{test}$ Accuracy (%)
+Table 6: Ablation study on different numbers of candidate documents and selected documents.
+$K_{candidate}$ $K_{train}$ $K_{test}$ Acc (%)
 5 1 1 57.90
@@
 30 5 1 61.62
-Effect of different vision information in our framework.
-In Tab.
-7, we explore the impact of different vision information under different settings.
-We directly finetune BLIP2 without using knowledge and we find that the model using LLM-aligned query features outperforms the model using visual contexts, improving the accuracy from 41.82% to 55.11%.
-Using both can bring the performance improvement a little.
-In our framework, using query features improves the VQA score from 52.93% to 62.83%.
-These results demonstrate that introducing LLM-aligned visual features can help the model understand the visual details better, which contributes to improving the model’s performance.
-Using both degrades the performance a little.
-Table 7. Ablation study on different vision information in our framework.
-Models Query features Context Accuracy (%)
-BLIP2 (fine-tuned) w/o knowledge $\checkmark$ 55.11
-$\checkmark$ 41.82
-$\checkmark$ $\checkmark$ 55.43
-Boter (Ours) $\checkmark$ 62.83
-$\checkmark$ 52.93
-$\checkmark$ $\checkmark$ 62.52
-Table 8. Ablation study on different documents selection in Answerer fine-tuning.
-Knowledge Selection Accuracy (%)
+Table 7: Ablation study on different documents selection in Answerer fine-tuning.
+Knowledge Selection Acc (%)
 Training Inference
@@
 Selector Selector 62.83
-Figure 3.
-Qualitative results on the test split of OK-VQA.
+Figure 2: Qualitative results on the test split of OK-VQA.
 We compared our method with a model that fine-tunes BLIP2 with knowledge ranked by DPR.
@@
 Tab.
-8 compares the Answerer fine-tuning with different document selection strategies.
+7 compares Answerer fine-tuning with different document selection strategies.
 The results show that our framework performs optimally when utilizing Selector in both Answerer training and inference.
 This is likely because the Selector provides more informative key knowledge documents and using both Selector ensures the consistency between the training domain and testing domain.
-4.4. Qualitative Analysis
+Performance of the knowledge retrieval.
+In tab.
+8, we evaluate our Selector in the knowledge retrieval task.
+Following previous methods Luo et al.
+(2021); Lin and Byrne (2022), we adopt pseudo relevance to measure if the retrieved document is relevant to the query due to the absence of ground-truth document.
+We use Recall to measure the performance of the the knowledge retrieval.
+From the results, we can see that our Selector improves the performance of DPR a lot.
+This means our Selector can retrieve more relevant knowledge documents which help answer questions.
+Compared to other retrievers, our Selector achieves the second best performance.
+Although FLMR outperforms our Selector in the knowledge retrieval, our framework achieves better accuracy in VQA (shown in Tab.
+1) with the same backbone.
+This indicates that the knowledge documents selected by Selector have better consistency with Answerer.
+Table 8: Retrieval performance on Google Search (GS).
+Retriever R@5 R@10
+VRR Luo et al. (2021) 80.4 88.55
+RA-VQA-FrDPR Lin and Byrne (2022) 81.25 88.51
+RA-VQA Lin and Byrne (2022) 82.84 89.00
+FLMR Lin et al. (2023) 89.32 94.02
+DPR Lin and Byrne (2022) 82.93 89.95
+Our Selector 88.66 93.56
+4.4 Qualitative Analysis
 In Fig.
-3, We present a case study comparing our method with a model that fine-tunes BLIP2 using knowledge ranked by DPR.
+2, We present a case study comparing our method with a model that fine-tunes BLIP2 using knowledge ranked by DPR.
 In the first case, top-ranked knowledge documents from DPR misguide the model, resulting in incorrect predictions.
@@
 These cases demonstrate our method’s ability to extract informative knowledge from retrieved documents to support accurate question answering.
-5. CONCLUSION
-In this paper, we propose Boter, a novel framework to address the task of knowledge-based VQA.
-We adopt the multimodal language model to construct two modules: a Selector and an Answerer.
-The Selector selects key knowledge from the retrieved knowledge documents and the Answerer reasons on the key knowledge to answer questions.
-Through a cycle training method, we ensure consistent improvement of both modules’ abilities in knowledge selection and question answering.
-Compared with state-of-the-art methods, Boter achieves better performance on a challenging open-domain knowledge-based VQA benchmark (OK-VQA) and we conduct a comprehensive analysis to evaluate the effectiveness of our method.
+5 Conclusion
+In this paper, we propose a novel framework that leverages the large visual-language model to construct two modules: (1) Selector for finding key retrieved knowledge and (2) Answerer for reasoning on the knowledge to predict answers.
+We design a self-bootstrap learning method to improve their abilities, where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector.
+Compared with state-of-the-art methods, our method achieves better performance on a challenging open-domain knowledge-based VQA benchmark (OK-VQA) and we conduct a comprehensive analysis to evaluate the effectiveness of our method.
+6 Acknowledgements
+This work was supported by the Artificial Intelligence-National Science and Technology Major Project (2023ZD0121200), the National Natural Science Foundation of China (U21B2043, 62102416), and the Key Research and Development Program of Jiangsu Province under Grant BE2023016-3.
+7 Limitations
+Although our framework can effectively select key knowledge documents for answering question, it is inevitable that the knowledge still contains noise.
+In some cases, the model itself can answer the question without external knowledge, introducing extra knowledge may affect the performance.
+In the future, we can explore to dynamically select required knowledge to help itself answer questions.
+In addition, there is a concern on the generalizability of the proposed method on other domains, especially when the initial DPR model can not retrieve gold standard context.
+In the future, we consider adopting a stronger multimodal retriever model to obtain more useful candidate knowledge documents, which enhances the generalizability of our framework.
+Appendix A Appendix
+Table 9: Performance comparison with state-of-the-art (SOTA) methods on the FVQA dataset.
+Method Acc-1
+Human 77.99
+UnifER Guo et al. (2022) 55.04
+FVQA Wang et al. (2017) 56.91
+ZS-VQA Chen et al. (2021) 58.27
+FVQA(Ensemble) Wang et al. (2017) 58.76
+MM-Reasoner(Ensemble) Khademi et al. (2023) 61.10
+Ours 63.3
+Table 10: Performance comparison with state-of-the-art (SOTA) methods on the A-OKVQA dataset.
+Method Direct Answer
+val test
+ClipCap Schwenk et al. (2022) 18.1 15.8
+Pythia Jiang et al. (2018) 25.2 21.9
+ViLBERT Lu et al. (2019) 30.6 25.9
+LXMERT Tan and Bansal (2019) 30.7 25.9
+KRISP Marino et al. (2021) 33.7 27.1
+GPV-2 Kamath et al. (2022) 48.6 40.7
+BLIP-2 T5-XL Li et al. (2023) 53.2 49.7
+PromptCap + GPT-3 Hu et al. (2022b) 56.3 59.6
+Ours 57.2 56.4
+A.1 Experiments on Other Datasets.
+We also evaluate our method on FVQA Fang et al.
+(2023) and A-OKVQA Schwenk et al.
+(2022) to demonstrate the effectiveness of our method.
+FVQA is a VQA dataset that mostly contains questions requiring external knowledge to answer, and provides supporting fact triplets alongside the image-question-answer triplets.
+A-OKVQA is an augmented successor of OK-VQA, containing 25K image-question pairs that require broader commonsense and world knowledge to answer.
+Due to A-OKVQA does not provide the knowledge source, we use Wikipedia Vrandečić and Krötzsch (2014) as the knowledge base.
+As shown in Tab.
+9, our method surpasses previous state-of-the-art methods, which demonstrates the effectiveness and generalization of our method.
+Tab.
+10 shows the comparative results on the challenging A-OKVQA dataset.
+Our method achieved competitive results, which demonstrates the effectiveness of our method.
+Table 11: Computational cost of our framework.
+$K_{candidate}$ Memory (GB) Running Time (sec./sample)
+10 21.3 1.0
+15 21.4 1.1
+30 22.7 1.3
+A.2 Evaluation of Computational Cost
+In Tab. 11, we show the computational cost of our framework using different knowledge candidates. As the number of candidate knowledge, the computational cost of our framework has a small increase.